OpenStudy (anonymous):

Toby is 5 feet tall. This is 4/5 the height of his father. How tall is Toby's father?

OpenStudy (anonymous):

Well if Toby is 5 feet tall then 5 = (4/5) of x where x = the height of his father.

OpenStudy (anonymous):

But if you multiply that you get 4

OpenStudy (anonymous):

No, you have to multiply each side by the reciprocal which in this case is 5/4. \[5/4 * 5 = 25/4 = 6.25\]

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

No prob yo =)